Discussion Question 1

You are a Clinical Director for Sanita Virtual Hospital. You just got your monthly financial report and saw that you were $8,260.00 above your budget of $202,480 on your staffing budget. You weren’t too concerned because you knew you had 111 more patient days than your expected budget of 930 patient days. You checked the staffing report and found that you paid 7458 hours against a budget of 7500.

What are the volume, efficiency and rate variances?

What should you be concerned about, if anything?

What should you do, if anything?
Answer #1

Answer:

Volume variances =(Standard Patient day-Actual) × Actual rate

= (903-1041)×202.44

=$3136.86 Advers

Actual rate = Actual exp./Actual patient day

=202480+8260/7458

=$28.26

Efficiency variance =(budgeted hour-Actual hours)×Std Rate

=(7500-7458)×27

=1134 Feverable

Std rate = Std oh /std hours

=202480÷7500

=26.99 ie. $27

Rate variance =(27-28.26)×7500

=9450 Adverse